OpenStudy (anonymous):

express as a single fraction in its simplest form 5/x - 3/(x-2)

OpenStudy (anonymous):

is this the problem: \(\Large \frac{5}{x}-\frac{3}{x-2} \) ???

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

get common denominators: the LCD is x(x-2) \(\Large \frac{5}{x}-\frac{3}{x-2} \) \(\Large \frac{5}{x}\cdot (\frac{x-2}{x-2})-\frac{3}{x-2}\cdot (\frac{x}{x}) \) can you take it from here?